OpenStudy (anonymous):

James Callahan has available a 10% alcohol solution and a 60% alcohol solution. Find how many liters of each solution he should mix to make 50 liters of a 40% alcohol solution?

OpenStudy (aaronq):

let x be the liters of 10% alcohol, and y be the liters of 60% alcohol, then: x+y=50 L 0.1 x +0.6y=0.4*50 L solve the system of equations.